40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 22 of 330  Not logged in ELOG logo
ID Date Author Type Category Subject
  15547   Sat Aug 29 20:07:48 2020 ranaUpdateElectronicsWFS characterization

I set up to do the WFS head modifications today, but I was shot down in flames due to a missing AC/DC adapter.

The Prologix GPIB-ethernet dongle needs +8-13 V to run. Some riff raff has removed the adapter and I was thunderstruck to see that it had not been returned.

I did the usual hunt around the lab looking for something with the right specs and connector. I found one that could do +9V and had the right connector, but it didn't light up the adapter so I put it back in black SP table.

I'll order a couple of these (5 ordered for delivery on Wednesday) in case there's a hot demand for the jack / plug combo that this one has. The setup is in the walkway, but I returned the AS table to the usual state and made sure the IMC is locking well.

  15546   Sat Aug 29 18:52:42 2020 ranaUpdateIOOIMC gain change

I lowered the (FAST) PZT gain on the IMC/FSS servo today.

I noticed that the MC locks looked unstable a lot of the day, and during lock the PCDRIVE channel is above 1 Vrms (which means the loop is oscillating, ttypically at the PZT/EOM crossover frequency).

I changed the default setting from 22 to 20 19 dB in the PSL Settings screen so the mcup script will use this for now. Feel free to revert if this turns out to be a Fluke (which you would think is a terrible name for a company, but...)

  15545   Fri Aug 28 23:33:38 2020 gautamUpdateBHDSome more hardware changes

Just a quick set of notes detailing changes so that there are no surprises, more details to follow.

  1. Trek driver has been temporarily placed on top of the KEPCO supply east of the OMC electronics rack. Cabling to it has been laid out as well. I turned both off so neither should be energized now.
  2. A new AI chassis (and associated cabling including the DAC SCSI cable and +/-24 V DC cable) has been installed in 1X2.
  3. To map the DAC range to what the Trek driver wants, I've configured the inverting summing amplifier with gain of 1/8. The offset voltage is set to 5V DC instead of 10V as intended, because the DAC can only drive +/-5 V when connected to a single ended receiving/sending unit.
  4. The LO delivery fiber was re-laid, and the interference between the IFO AS beam and LO beams were restored.

I briefly tried some LO PZT mirror dithering tonight, but didn't see the signal. Needs more troubleshooting.

  15544   Fri Aug 28 11:41:23 2020 gautamBureaucracysafetyCrane inspection 2020

Mr Fred Goodbar of Konacrane was in the lab 830am-1130am today. All three cranes in the VEA were inspected, loaded with 450lb test weights, and declared in good working condition and safe to use.

  1. Apparently, the clackity noise heard when running the crane at the south end is a known problem - the crane was opened up and inspected sometime in the past, and no obvious cause was found. This is not expected to affect the usability of the crane.
  2. The travel speed of the cranes is slow - but this is apparently intentional, on the request of Steve V.

The interferometer subsystems appear normal after the inspection. 

  15543   Wed Aug 26 22:49:47 2020 gautamUpdateElectronicsCheckout of Trek Model 603

I unboxed the Trek amplifier today, and performed some basic tests of the functionality. It seems to work as advertised. However, we may have not specified the correct specifications - the model seems to be configured to drive a bipolar output of +/- 125 V DC, whereas for PZT driving applications, we would typically want a unipolar drive signal. From reading the manual, it appears to me that we cannot configure the unit to output 0-250V DC, which is what we'd want for general PZT driving applications. I will contact them to find out more. 

The tests were done using the handheld precision voltage source for now. I drove the input between 0 to +5 V and saw an output voltage (at DC) of 0-250 V. This is consistent with the voltage gain being 50V/V as is stated in the manual, but how am I able to get 250 V DC output even though the bipolar configuration is supposed to be +/- 125 V? On the negative side, I am able to see 50V/V gain from 0 to -1 V DC. At which point making the input voltage more negative does nothing to the output. The unit is supposed to accept a bipolar input of +/- 10 V DC or AC, so I'm pretty sure I'm not doing anything crazy here...

Update:

Okay based on the markings on the rear panel, the unit is in fact configured for unipolar output. What this means is we will have to map the +/- 10 V DC output from the DAC to 0-5 V DC. Probably, I will stick to 0-2.5 V DC for a start, to not exceed 125 V DC to the PI PZT. I'm not sure what the damage spec is for that. The Noliac PZT I think can do 250 V DC no problem. Good thing I have the inverting summing amplifier coming in tomorrow...

Attachment 1: IMG_8951.JPG
IMG_8951.JPG
  15542   Wed Aug 26 16:12:25 2020 gautamUpdateElectronicsTest mass coil current requirements

Attachment #1 is a summary of the current to each coil on the suspensions. The situation is actually a little worse than I remembered - several coils are currently drawing in excess of 10mA. However, most of this is due to a YAW correction, which can be fixed somewhat more easily than a PIT correction. So I think the circuit with a gain of 31 for an input range of +/-10 V, which gives us the ability to drive ~12mA per coil through a 25kohm series resistor, will still provide sufficient actuation range. As far as the HV supplies go, we will want something that can do +/- 350 V. Then the current to the coils will at most be ~50 mA per optic. The feedback path will require roughly the same current. The quiescent draw of each PA95 is ~10mA. So per SOS suspension, we will need ~150mA.

If it turns out that we need to get more current through the 25kohm series resistance, we may have to raise the voltage gain of the circuit. Reducing the series resistance isn't a good option as the whole point of the circuit is to be limited by the Johnson noise of the series resistance. Looking at these numbers, the only suspension on which we would be able to plug in a HV coil driver as is (without a vent to correct for YAW misalignment) is ITMY.


Update 2 Sep 2020 2100: I confirmed today that the number reported in the EPICS channel, and the voltage across the series resistor, do indeed match up. The test was done on the MC3 coil driver as it was exposed and I didn't need to disable any suspensions. I used a Fluke DMM to measure the voltage across the resistor. So there is no sneaky factor of 2 as far as the Acromag DACs are concerned (unlike the General Standards DAC).

Attachment 1: coilCurrents.png
coilCurrents.png
  15541   Wed Aug 26 15:48:31 2020 gautamUpdateVACControl screen left open on vacuum workstation

I found that the control MEDM screen was left open on the c1vac workstation. This should be closed every time you leave the workstation, to avoid accidental button pressing and such.

The network outage meant that the EPICS data from the pressure gauges wasn't recorded until I reset everything ~noon. So there isn't really a plot of the outgassing/leak rate. But the pressure rose to ~2e-4 torr, over ~4 hours. The pumpdown back to nominal pressure (9e-6 torr) took ~30 minutes.

  15540   Wed Aug 26 00:52:55 2020 gautamUpdateBHDBHD activities

Listing some talking points from the last week of activity here.

  1. LO delivery fiber cable may be damaged.
    • The throughput itself doesn't suggest any problems, I get almost all the light I put in out the other end.
    • However, even when I slightly move the fiber, I see huge amplitude fluctuations in the DCPD readouts. This shouldn't be the case, particularly if the light is well matched to one of the special axes of the PM fiber. I checked with a PBS at the output that this is indeed the case, so something else must be funky?
    • In any case, I don't think it's a great idea to use this 70m long fiber for bringing the light from the PSL table to the adjacent AP table. Chub has ordered a 10m patch cable.
    • I was a bit too hasty this morning, thinking we had a patch cable in hand, and so I removed the fiber from the AP table. So right now, the LO beam doesn't make it to the BHD setup. Depending on the lead time for the new patch cable, I may or may not resurrect this old setup.
    • I have also located some foam and rigid plastic tubing which I think will help in isolating the fiber from environmental length(phase) modulation due to acoustic pickup.
  2. BHD commissioning activities
    • Basically, I've been trying to use the Single Bounce ITM reflection/ Michelson / PRMI with carrier locked to get some intuition about the BHD setup. These states are easily prepared, and much easier to understand than the full IFO for these first attempts.
    • One concern I have is the angular stability (or lack thereof). When the PRMI is locked, the DC light level on each DCPD fluctuates between ~0 (which is what it should be), up to ~30 cts (~85uW).
    • Using the empirically determined attenuation factor between the DCPDs and the dark port of the beamsplitter, I estimate the power can be as high as 20mW. This is a huge number, considering the input to the interferometer is ~800mW. I assume that all the light is at the carrier frequency, since the PRC should reject all the sideband light in this configuration. In any case, the total amount of sideband light is ~20mW, and the carrier stays resonant in the PRC even when there are these large ASDC excursions, so I think it's a reasonable assumption that the light is at the carrier frequency. Moreover, looking at the camera, one can see a clear TEM10/01 profile, indicative of imperfect destructive interference at the beamsplitter due to beam axis misalignment.
    • The effect of such excursions on the BHD readout hasn't yet been quantified (by me at least), but I think it may be hampering my attempts to dither the homodyne phase to estimate the LO phase noise.
  3. High voltage coil driver project - see thread for updates.
  4. Trek HV driver has arrived.
    • I haven't opened the box yet, but basically, what this means is that I can dither the mirror intended for homodyne phase control in a reasonable way.
    • Previously, I was using the OMC HV driver to drive the PZTs - but this dither signal path has a 2kHz high pass filter (since the OMC length dither is a kHz dither). I didn't want to futz around with the electronics, particularly since the unit was verified to be working.
    • So the plan now would be to drive the input of the Trek with a DAC output (an appropriate AI chassis has been prepared to interface with the CDS system).
    • Hopefully, there's enough DAC dynamic range to dither the PZT and also do the homodyne phase locking using a single channel. Else, we'd need to use two channels and install a summing amplifier.
    • We definitely need more high-voltage amplifiers/supplies in the lab:
      • Any Thorlabs HV drivers we can recover? 
      • Eventually, we will need HV for coil drivers, OMC PZTs, steering PZTs, homodyne phase control PZT. 
  5. PMC bases have arrived.
    • Joe Benson from the machine shop informed me today afternoon that the bases were ready for pickup.
    • We have 3 bases in hand now. The finish isn't the greatest in the world, but I think it'll work. You can see some photos here.
    • I will hold off on putting this together while I work on the basic airBHD commissioning tests. We can install the PMCs later.
  6. AS port WFS project
    • We now have in hand almost all the components for stuffing the ISC whitening and LSC demod boards.
    • Rich, Chub, Luis and I had a call on Monday. The advise from Rich/Luis was:
      • Choose an inductance that has Z~100 ohms at the frequency of interest, for the resonant transimpedance part.
      • Choose a capacitance that gives the appropriate resonant frequency.
      • Don't stuff more notches than you need - start with just a 2f notch (so 110 MHz for us), and make sure to place the highest frequency notch closest to the photodiode.
      • Rich also suggested looking at the optical signal with a non-optimized head, get an idea of what the field content is, and then tune the circuit as necessary. There are obviously going to be many issues that only become apparent once we do such a test.
    • The aLIGO modulation frequencies are only 20% different from the 40m modulation frequencies. So I thought it is best if for our first pass, we stick to the inductance values used in the aLIGO circuits (same footprint, known part etc etc). Then, we will change the capacitance so that we have a tuning range that is centered our modulation frequencies.
    • The parts have been ordered.
  7. ISS project
    • Half of the LO light on the BHD breadboard is diverted for the purpose of sensing the LO intensity noise, for eventual stabilization. Right now, it is just getting dumped.
    • A PD head has been located. It has a minimalist 1kohm transimpedance amplifier circuit integrated into the head.
    • Our AOM driver has an input range of 0-1V DC. We want to map the servo output of +/-10V DC (or +/-4V DC if we use an SR560 based servo for a first pass) to this range.
    • I wanted to do this for once in a non-hacky way so I drew up a circuit that I think will serve the purpose. It has been fabricated and will be tested on the bench in a couple of days.
    • Once I get a feel for what the signal content is, I will also draw up a interface board to the PD head that (i) supplies the reverse bias voltage and +/-15 V DC to the PD head and (ii) applies some appropriate HPF action and provides a DC monitor as well.
  8. Summary pages are dead.
  9. General lab cleanup
    • I moved all the PPE from the foyer area into the designated cabinets along the east arm.
    • Did some basic cleanup of the lab in preparation for crane inspection. Walkways are clear.
    • I de-cluttered the office area a bit, but today I received ~10 packages from Digikey/FrontPanelExpress etc. So, in fact, it got even more cluttered. Entropy will go down once we ship these off to screaming circuits for stuffing the PCBs.
  15539   Tue Aug 25 05:51:29 2020 YehonathanUpdateBHDMonte Carlo Simulations

I re-plotted the MCMC results as semi-transparent lines so that probable lines stick out.

This also reveals what is behind the extreme sparsity in the aLIGO simulation results (In the previous post).

There seem to be some bi-stability/phase transition/whatever in the aLIGO simulation. The aLIGO transfer functions are very sensitive to one or more of the DOFs. Not sure which yet.

Attachment 1: MICH_AplusMCMC.pdf
MICH_AplusMCMC.pdf
Attachment 2: PRCL_AplusMCMC.pdf
PRCL_AplusMCMC.pdf
Attachment 3: SRCL_AplusMCMC(1).pdf
SRCL_AplusMCMC(1).pdf
Attachment 4: OMC_Diff_AplusMCMC.pdf
OMC_Diff_AplusMCMC.pdf
Attachment 5: OMC_Comm_AplusMCMC.pdf
OMC_Comm_AplusMCMC.pdf
Attachment 6: OMC_Angle_Yaw_AplusMCMC.pdf
OMC_Angle_Yaw_AplusMCMC.pdf
Attachment 7: OMC_Angle_Pitch_AplusMCMC.pdf
OMC_Angle_Pitch_AplusMCMC.pdf
Attachment 8: Main_Laser_RIN_AplusMCMC.pdf
Main_Laser_RIN_AplusMCMC.pdf
  15538   Mon Aug 24 11:25:07 2020 JonUpdateVACUPS installation

I'm leaving the lab shortly. We're not ready to switch over the vac equipment to the new UPS units yet.

The 120V UPS is now running and interfaced to c1vac via a USB cable. The unofficial tripplite python package is able to detect and connect to the unit, but then read queries fail with "OS Error: No data received." The firmware has a different version number from what the developers say is known to be supported.

The 230V UPS is actually not correctly installed. For input power, it has a general type C14 connector which is currently plugged into a 120V power strip. However this unit has to be powered from a 230V outlet. We'll have to identify and buy the correct adapter cable.

With the 120V unit now connected, I can continue to work on interfacing it with python remotely. The next implementation I'm going to try is item #2 of this plan [ELOG 15446].

Quote:

I'm in the lab this morning to interface the two new UPS units with the digital controls system. Will be out by lunchtime. The disruptions to the vac system should be very brief this time.

  15537   Mon Aug 24 08:13:56 2020 JonUpdateVACUPS installation

I'm in the lab this morning to interface the two new UPS units with the digital controls system. Will be out by lunchtime. The disruptions to the vac system should be very brief this time.

  15536   Sun Aug 23 23:36:58 2020 gautamUpdateElectronicsFirst look at HV coil driver

Summary:

A more careful analysis has revealed some stability problems. I see oscillations at frequencies ranging from ~600kHz to ~1.5 MHz, depending on the voltage output requested, of ~2 V pp at the high-voltage output in a variety of different conditions (see details). My best guess for why this is happening is insufficient phase margin in the open-loop gain of the PA95 high voltage amplification stage, which causes oscillations to show up in the closed loop. I think we can fix the problem by using a larger compensation capacitor, but if anyone has a better suggestion, I'm happy to consider it

Details:

The changes I wanted to make to the measurement posted earlier in this thread were: (i) to measure the noise with a load resistor of 20 ohms (~OSEM coil resistance) connected, instead of the unloaded config previously used, and (ii) measure the voltage noise on the circuit side (= TP5 on the schematic) with some high voltage output being requested. The point was to simulate conditions closer to what this board will eventually be used in, when it has to meet the requirement of <1pA/rtHz current noise at 100 Hz. The voltage divider formed by the 25 kohm series resistor and the 20 ohm OSEM coil simulated resistance makes it hopeless to measure this level of voltage noise using the SR785. On the other hand, the high voltage would destroy the SR785 (rated for 30 V max input). So I made a little Pomona box to alllow me to do this measurement, see Attachment #1. Its transfer function was measured, and I confirmed that the DC high voltage was indeed blocked (using a Fluke DMM) and that the output of this box never exceeded ~1V, as dictated by the pair of diodes - all seemed okay .

Next, I wanted to measure the voltage noise with ~10mA current flowing through the output path - I don't expect to require more than this amount of current for our test masses. However, I noticed some strange features in the spectrum (viewed continuously on the SR785 using exponential averaging setting). Closer investigation using an oscilloscope revealed:

  1. 600kHz to 1 MHz oscillations visible, depending on output voltage.
  2. The oscillations vanish if I drive output above +30 V DC (so input voltage > 1 V).
  3. The oscillations seem to be always present when the output voltage is negative.
  4. No evidence of this offset if circuit is unloaded and voltage across 25k resistor is monitored. But they do show up on scope if connected to circuit side even in this unloaded config.

Some literature review suggested that the capacitor in the feedback path, C4 on the schematic, could be causing problems. Specifically, I think that having that capacitor in the feeddback path necessitates the use of a larger compensation capacitor than the nominal 33pF value (which itself is higher than the 4.7pF recommended on the datasheet, based on experience of the ESD driver circuit which this is based on, oscillations were seen there too but the topology is a bit different). As a first test of this idea, I removed the feedback capacitor, C4 - this seemed to do the trick, the oscillations vanished and I was able to drive the output between the high voltage supply rails. However, we cannot operate in this configuration because we need to roll off the noise gain for the input voltage noise of the PA95 (~6 nV/rtHz at 100 Hz will become ~200 nV/rtHz, which I confirmed using the SR785). Using a passive RC filter at the output of the PA95 (a.k.a. a "snubber" network) is not an option because we need to sum in the fast actuation path voltage at the output of the 25 kohm resistor.

Some modeling confirms this hypothesis, see Attachment #2.  The quantity plotted is the open-loop gain of the PA95 portion of the circuit. If the phase is 0 degrees, then the system goes unstable.

So my plan is to get some 470pF capacitors and test this idea out, unless anyone has better suggestions? I guess usually the OpAmps are compensated to be unconditionally stable, but in this case maybe the power op-amp is more volatile?

Quote:

Need to think more about how to better characterize this noise. An estimate of the required actuation can be found here.

Attachment 1: IMG_5379.JPG
IMG_5379.JPG
Attachment 2: stabilityCriterion.pdf
stabilityCriterion.pdf
  15535   Fri Aug 21 15:27:00 2020 gautamUpdateBHDBetter BHD mode-matching

Summary:

The mode-matching between the LO and AS beams is now ~50%. This isn't probably my most average mode-matching in the lab, but I think it's sufficient to start doing some other characterization and we can try squeezing out hopefully another 20-30% by putting the lenses on translation stages, tweaking alignment etc.

Details:

The main change was to increase the optical path length of the IFO AS path, see Attachment #1. This gave me some more room to put a lens and translate it.

  • The LO path uses two lenses, f=200mm and f=100mm to focus the collimator output beam, which is supposedly ~1200um diameter, to something like 400um diameter (measured using beam profiler but not very precisely).
  • This beam is  fairly well collimated, and the beam size is close to what the PMC cavity will want, I opted not to tweak this too much more.
  • For the AS beam, the single bounce reflection from ITMY was used for alignment work.
  • There is a 2" f=600mm lens upstream (not seen in Attachment #1). This supposedly makes a beam with waist ~80um, but I couldn't numerically find a good solution numerically if this assumption is true, so I decided to do the mode-matching empirically.
  • A single f=150mm lens got me a beam that seemed pretty well collimated, and roughly the same size as the LO beam, so I opted to push ahead with that. Later, I measured with the beam profiler that the beam is ~600um in diameter, so the beam isn't very well matched to the LO spot size, but I decided to push ahead nevertheless.
  • Patient alignment work enabled me to see interference fringes.
    • Note that the ITM reflection registers 30 cts (~80 uW). Assuming 800mW transmission through the IMC, I would have expected more like 800mW * 5.637% * 50% * 98.6% * 50% * 10% * 30% * 50% * 50% = 80uW, so this is reasonable I guess. The chain of numbers corresponds to T_PRM * T_BS * R_ITM * R_BS * T_SRM * T_vac_OMC_pickoff * R_in_air_BS * R_homodyneBS.
    • The IFO AS beam appears rather elliptical to the eye (and also on the beam profiler). It already looks like this coming out of the vacuum so not much we can do about it right now I guess. By slightly rotating the f=150mm focusing lens so that the beam going through it at ~10 degrees instead of normal incidence, I was able to get a more circular beam as measured using the beam profiler.
    • With the AS beam blocked, the LO beam registers 240 cts on each DCPD (~0.7 mW). 
    • The expected fringe should then be (sqrt(240) + sqrt(30))^2 - (sqrt(240) - sqrt(30))^2 ~ 440 cts pp.
    • The best alignment I could get is ~200 cts pp, see Attachment #2.

Next steps:

Try the PRMI experiments again, now that I have some confidence that the beams are actually interfering.

See Attachment #3 for the updated spectra - the configuration is PRMI locked with carrier resonant and the homodyne phase is uncontrolled. There is now much better clearance between the electronics noise and the MICH signal as measured in the DCPDs. The "LO only" trace is measured with the PSL shutter closed, so the laser frequency isn't slaved to the IMC length. I wonder why the RIN (seen in the SUM channel) is different whether the laser is locked to the IMC or not? The LO pickoff is before the IMC.

Attachment 1: IMG_7548.JPG
IMG_7548.JPG
Attachment 2: BHD_MM.png
BHD_MM.png
Attachment 3: PRMI_DCPDs.pdf
PRMI_DCPDs.pdf
  15534   Thu Aug 20 00:21:51 2020 gautamUpdateElectronicsFirst look at HV coil driver

Summary:

A single channel of this board was stuffed (and other channels partially populated). The basic tests passed, and nothing exploded! Even though this is a laughably simple circuit, it's nice that it works.

HV power supplies:

A pair of unused KEPCO BHK300-130 switching power supplies that I found in the lab were used for this test. I pulled the programmable cards out at the rear, and shorted the positive output of one unit to the negative of the other (with both shorted to the supply grounds as well), thereby creating a bipolar supply from these unipolar models. For the purposes of this test, I set the voltage and current limits to 100V DC, 10mA respectively. I didn't ramp up the supply voltage to the rated 300 V maximum. The setup is shown in Attachment #1.

Tests:

  1. With the input to the channel shorted to ground, I confirmed with a DMM that the output was (nearly) zero (there was an offset of ~40mV but I think this is okay).
  2. Used the calibrated voltage source, and applied +/- 3 V in steps of ~0.5 V, while monitoring the output with a DMM. Confirmed the output swing of ~ +/-90 V, which is what is expected, since the design voltage gain of this circuit is 31.
  3. Drove a 0.1 Hz, 500mVpp sine wave at the input while monitoring the output and the Vmon testpoints, see Attachment #2. Note the phasing between input and output, and also the fact that the gain is slightly lower than the expected gain of 31, because there are three poles at ~0.7 Hz, which already start showing some influence on the transfer function at 0.1 Hz.
  4. Noise measurement 
    • The whole point of this circuit is to realize sub 1pA/rtHz current noise to the coil, when it is connected.
    • For this test, no load was connected (i.e. voltage noise was measured at the output of the 25 kohm resistor), and the input was shorted to ground so that the DC value of the output was close to 0 (the idea was to not overload the SR560/SR785 with high voltage).
    • An SR560 preamp with gain x50 (DC coupled) was used to preamplify the signal. This was the maximum gain that could be used with the unit DC coupled, due to the small DC offset. I opted to keep the DC coupling to get a look at the low frequency noise as well, but in hindsight, maybe I should have used AC coupling as we only care about the current noise at ~100 Hz.
    • See Attachment #3 for results. The measurement is close to the model above ~100 Hz

Need to think more about how to better characterize this noise. An estimate of the required actuation can be found here.

Attachment 1: IMG_8724.JPG
IMG_8724.JPG
Attachment 2: timeDomain.pdf
timeDomain.pdf
Attachment 3: HVampNoise.pdf
HVampNoise.pdf
  15533   Tue Aug 18 13:55:23 2020 ranaUpdateALSWhitening and ALS noise

No, there should be no unscheduled visits from any inspector, marshal, tech, or vendor. They all have to be escorted or they don't get in. If they have a problem with that, please give them my cell #.

 

For the ALS, in addition to the beat note spectrum, I think we need to know the loop gain use to feedback to the ETM to determine the true cavity length fluctuation. w/o ALS, the noise would be only due to the seismic noise, OSEM damping noise, and the IR-PDH residual. Those are all suppressed by the ALS loop, but then the ALS loop puts its sensing noise onto the cavity. So, if I'm thinking about this right, the ALS beat noise > 200 Hz doesn't matter so much to the CARM RMS. So the whitening seems to be doing good in the right spot, but we would like to have another boost in the green PDH to up the gain below ~300 Hz?

  15532   Mon Aug 17 23:41:50 2020 gautamUpdateBHDWhitening and air BHD dark noise

Summary:

With the chosen transimpedance of 300 ohms, in order to be able to see the shot noise of 10 mW of light in the digitized data streams, we'd need all 3 stages of whitening. If we want to be shot noise limited with 1 mW of LO light, we'd need to increase said transimpedance I think.

Details:

The measurements were taken with

  1. No light incident on the DCPDs.
  2. The flat whitening gain was set to 0 dB.
  3. Whitening engaged sequentially, stage by stage, shown as (Blue, Red, Orange and Green) curves corresponding to (0, 1, 2, 3) stages of whitening.

Of course, it's unlikely we're going to be shot noise limited for any configuration in the short run. But this was also a test of 

  1. My soldering.
  2. Change of whitening corner frequencies.
  3. Test of the overall whitening board assembly.

All 3 tests passed.

Attachment 1: BHD_whitening.pdf
BHD_whitening.pdf
  15531   Mon Aug 17 23:36:10 2020 gautamUpdateALSWhitening and ALS noise

finally managed to install a differential-receiving whitening board in 1Y2 - 4 channels are available at the moment. As I claimed, one stage of 15:150 Hz z:p whitening does improve the ALS noise a little, see Attachment #1. While the RMS (from 1kHz-0.5 Hz) does go down by ~10 Hz, this isn't really going to make any dramatic improvement to the 40m lock acquisiton. Now we're really sitting on the unsuppressed EX laser noise above ~30 Hz. This measurement was taken with the arm cavities locked with POX/POY, and end lasers locked to the arm cavities with uPDH boxes as usual. This was just a test to confirm my suspicion, the whitening board is to be used for the air BHD channels, but when we get a few more stuffed, we can install it for the ALS channels too.

Attachment 1: ALSimprovement.pdf
ALSimprovement.pdf
  15530   Mon Aug 17 21:24:43 2020 gautamUpdateGeneralFire extinguisher inspection

A technician came to the lab today at ~4pm. He entered the VEA (with booties and googles), and also the clean and bake lab. The whole procedure lasted ~10 minutes. I did not follow him around, but was available in the control room throughout the process. I think the whole episode went without incident.

BTW, this guy didn't ring the doorbell, I just happened to be here when he came by. I don't know if this is usual practise - are we happy with the technicians entering the VEA and/or clean and bake labs without supervision? AFAIK, this wasn't scheduled.

  15529   Mon Aug 17 15:18:26 2020 gautamUpdateEquipment loanBeam Profiler + peripherals --> 40m

Gabriele left the DataRay beam profiler + peripherals (see Attachment #1) in his office. I picked them up just now and brought them over to the 40m.

Attachment 1: IMG_8719.JPG
IMG_8719.JPG
  15528   Sat Aug 15 15:12:22 2020 JonConfigurationVACOverhaul of small turbo pump interlocks

Summary

Yesterday I completed the switchover of small turbo pump interlocks as proposed in ELOG 15499. This overhaul altogether eliminates the dependency on RS232 readbacks, which had become unreliable (glitchy) in both controllers. In their place, the V4(5) valve-close interlocks are now predicated on an analog controller output whose voltage goes high when the rotation speed is >= 80% of the nominal setpoint. The critical speed is 52.8 krpm for TP2 and 40 krpm for TP3. There already exist hardware interlocks of V4(5) using the same signals, which I have also tested.

Interlock signal

Unlike the TP1 controller, which exposes simple relays whose open/closed states are sensed by Acromags, what the TP2(3) controllers output is an energized 24V signal for controlling such a relay (output circuit pictured below). I hadn't appreciated this difference and it cost me time yesterday. The ultimate solution was to route the signals through a set of new 24V Phoenix Contact relays installed inside the Acromag chassis. However, this required removing the chassis from the rack and bringing it to the electronics bench (rather than doing the work in situ, as I had planned). The relays are mounted to the second DIN rail opposite the Acromags. Each TP2(3) signal controls the state of a relay, which in turn is sensed using an Acromag XT1111.

Signal routing

The TP2(3) "normal-speed" signals are already in use by hardware interlocks of V4(5). Each signal is routed into the main AC relay box, where it controls an "interrupter" relay through which the Acromag control signal for the main V4(5) relay is passed. These signals are now shared with the digital controls system using a passive DB15 Y-splitter. The signal routing is shown below.

Interlock conditions

The new turbo-pump-related interlock conditions and their channel predicates are listed below. The full up-to-date channel list and wiring assignments for c1vac are maintained here.

Channel Type New? Interlock-triggering condition
C1:Vac-TP1_norm BI No Rotation speed < 90% nominal setpoint (29 krpm)
C1:Vac-TP1_fail BI No Critical fault occurrence
C1:Vac-TP1_current AI No Current draw > 4 A
C1:Vac-TP2_norm BI Yes Rotation speed < 80% nominal setpoint (52.8 krpm)
C1:Vac-TP3_norm BI Yes Rotation speed < 80% nominal setpoint (40 krpm)

There are two new channels, both of which provide a binary indication of whether the pump speed is outside its nominal range. I did not have enough 24V relays to also add the C1:Vac-TP2(3)_fail channels listed in ELOG 15499. However, these signals are redundant with the existing interlocks, and the existing serial "Status" readback will already print failure messages to the MEDM screens. All of the TP2(3) serial readback channels remain, which monitor voltage, current, operational status, and temperature. The pump on/off and low-speed mode on/off controls remain implemented with serial signals as well.

The new analog readbacks have been added to the MEDM controls screens, circled below:

Other incidental repairs

  • I replaced the (dead) LED monitor at the vac controls console. In the process of finding a replacement, I came across another dead spare monitor as well. Both have been labeled "DEAD" and moved to Jordan's desk for disposal.
  • I found the current TP3 Varian V70D controller to be just as glitchy in the analog outputs as well. That likely indicates there is a problem with the microprocessor itself, not just the serial communications card as I thought might be the case. I replaced the controller with the spare unit which was mounted right next to it in the rack [ELOG 13143]. The new unit has not glitched since the time I installed it around 10 pm last night.
Attachment 1: small_tp_signal_routing.png
small_tp_signal_routing.png
Attachment 3: small_tp_signal_routing.png
small_tp_signal_routing.png
Attachment 4: medm_screen.png
medm_screen.png
  15527   Sat Aug 15 02:02:13 2020 JonConfigurationVACVacuum repairs today

Vacuum work is completed. The TP2 and TP3 interlocks have been overhauled as proposed in ELOG 15499 and seem to be performing reliably. We're now back in the nominal system state, with TP2 again backing for TP1 and TP3 pumping the annuli. I'll post the full implementation details in the morning.

I did not get to setting up the new UPS units. That will have to be scheduled for another day.

Quote:

The vac system is going down now for planned repairs [ELOG 15499]. It will likely take most of the day. Will advise when it's back up.

  15526   Fri Aug 14 10:10:56 2020 JonConfigurationVACVacuum repairs today

The vac system is going down now for planned repairs [ELOG 15499]. It will likely take most of the day. Will advise when it's back up.

  15525   Fri Aug 14 10:03:37 2020 JonUpdateCDSTiming distribution slot availability

That's great news we won't have to worry about a new timing fanout for the two new machines, c1bhd and c1sus2. And there's no plan to change Dolphin IPC drivers. The plan is only to install the same (older) version of the driver on the two new machines and plug into free slots in the existing switch.

Quote:

The new dolphin eventually helps us. But the installation is an invasive change to the existing system and should be done at the installation stage of the 40m BHD.

  15524   Fri Aug 14 00:01:55 2020 gautamUpdateCDSBHD / OMC model channels now added to autoburt

I added the EPCIS channels for the c1omc model (gains, matrix elements etc) to the autoburt such that we have a record of these, since we expect these models to be running somewhat regularly now, and I also expect many CDS crashes.

  15523   Thu Aug 13 18:10:22 2020 gautamUpdateGeneralPower outage

There was a power outage ~30 mins ago that knocked out CDS, PSL etc. The lights in the office area also flickered briefly. Working on recovery now. The elog was also down (since nodus presumably rebooted), I restarted the service just now. Vacuum status seems okay, even though the status string reads "Unrecognized".

The recovery was complete at 1830 local time. Curiously, the EX NPRO and the doubling oven temp controllers stayed on, usually they are taken out as well. Also, all the slow machines and associated Acromag crates survived. I guess the interruption was so fleeting that some devices survived.

The control room workstation, zita, which is responsible for the IFO status StripTool display on the large TV screen, has some display driver issues I think - it crashed twice when I tried to change the default display arrangement (large TV + small monitor). It also wants to update to Ubuntu 18.04 LTS, but I decided not to for the time being (it is running Ubuntu 16.04 LTS). Anyways, after a couple of power cycles, the wall StripTools are up once again.

  15522   Thu Aug 13 13:35:13 2020 KojiUpdateCDSTiming distribution slot availability

The new dolphin eventually helps us. But the installation is an invasive change to the existing system and should be done at the installation stage of the 40m BHD.

  15521   Thu Aug 13 11:30:19 2020 gautamUpdateCDSTiming distribution slot availability

That's great. I wonder if we can also get away with not adding new Dolphin infrastructure. I'd really like to avoid changing any IPC drivers.

Quote:

I believe we will use two new chassis at most. We'll replace c1ioo from Sun to Supermicro, but we recycle the existing timing system.

  15520   Wed Aug 12 20:16:52 2020 KojiUpdateElectronicsPhotodiode inventory

When I tested Q3000 for aLIGO, the failure rate was pretty high. Let's get 10pcs.

  15519   Wed Aug 12 20:15:42 2020 KojiUpdateElectronicsNumber of the beast

Grrr. Let's repair the unit. Let's get a help from Chub & Jordan.

Do you have a second unit in the lab to survive for a while?

  15518   Wed Aug 12 20:14:06 2020 KojiUpdateCDSTiming distribution slot availability

I believe we will use two new chassis at most. We'll replace c1ioo from Sun to Supermicro, but we recycle the existing timing system.

  15517   Wed Aug 12 18:08:54 2020 gautamUpdateElectronicsNumber of the beast

The "source" output of the SR785 has a DC offset of -6.66 V. I couldn't make this up.

Upshot is, this SR785 is basically not usable for TF measurements. I was using the unit to characterize the newly stuffed ISC whitening board. The initial set of measurements were sensible, and at some point, I started getting garbage data. Unclear what the cause of this is. AFAIK, we don't have any knob to tune the offset - adjusting the "offset" in the source menu, I can change the level of the offset, but only by ~1 V even if I apply an offset of 10 V. I also tried connecting the ground connection on the rear of the SR785 to the bench power supply ground, no change.

Do we have to send this in for repair?

Attachment 1: IMG_8710.JPG
IMG_8710.JPG
  15516   Wed Aug 12 17:42:58 2020 gautamUpdateElectronicsPhotodiode inventory

See Attachments #1 and #2. We don't have any Q3000 QPDs in hand, at least not in the photodiode box stored in the clean optics cabinet at the south end. I also checked a cabinet along the east arm where we store some photodiodes - but didn't find any there either. The only QPDs we have in hand are the YAG-444-4AH, which I believe is what is used in the iLIGO WFS heads.

So how many do we want to get?

Attachment 1: IMG_8709.JPG
IMG_8709.JPG
Attachment 2: IMG_8708.JPG
IMG_8708.JPG
  15515   Wed Aug 12 17:36:42 2020 gautamUpdateCDSTiming distribution slot availability

See Attachment #1. J8 was connected to a "LASTI timing slave" sitting in the rack that Chiara lives in - we don't use this for anything and I confirmed that there was no effect on the RTCDS when I pulled that fiber out. The LASTI timing slave also had a blinky that was blinking when the fiber was plugged in - which I take to believe that the slot works. 

Can we get away with just using these two available slots, J8 and J13? Do we really need three new expansion chassis?

Attachment 1: IMG_8706.JPG
IMG_8706.JPG
  15514   Tue Aug 11 23:20:29 2020 gautamUpdateBHDSome first tests with air BHD setup

Some tests done today:

All of these tests were done with the PRMI locked with carrier resonant in the recycling cavity (i.e. sidebands rejected to REFL port). I then actuated the BS length DOF with a sine wave at 311.1 Hz, 40 cts amplitude (corresponding to ~8 pm of peak-to-peak displacement).

  1. Attempt to balance the DCPDs
    • I tried to tune the digital gains of the two DCPDs so as to minimize the appearance of this line in the SUM channel
    • but no matter how I tuned the gains, I couldn't make the line in the SUM channel disappear entirely - in fact, the best I could do was to make the line height in SUM and NULL channels (yes I recognize the poor channel name choice, I'll change "NULL" to "DIFF" at the next model recompile) the same. See Attachment #1.
    • The lobes around the main peak are indicative of some scattering?
    • Attachment #2 shows a wider frequency range. The homodyne phase isn't controlled, so the "NULL" channel is not necessarily measuring the correct quadrature to be sensing MICH motion.
    • I think I can back out something about the contrast defect from this fact, but I need to go back to some modeling.
  2. A simple test of the homodyne phase actuator
    • I wanted to check that this PI S320 piezo actually allows me to actuate the optical path length of the local oscillator.
    • I'm using the OMC HV driver to drive said PZT - so there are two DAC channels available, one to dither the optic and one to apply a control signal. I think mainly this is to avoid using up DAC range for the dither signal, the overall dynamic range is still limited by the HV supply.
    • I can't find the maximum voltage that can be applied on the datasheet - so conservatively, I limited the HV output to saturate at 100 V DC, as this is the maximum for the S330 piezos used for green steering, for which there is a manual.
    • The S320 manual does say the full stroke of each PZT element is 10 um - so the actuation coefficient is ~100 nm/V. I then drove this actuator with a sine wave of 500 cts amplitude, at 314.1 Hz (corresponding to 15 nm of motion). With only the LO beam incident on the PDs, I saw no signal in either DCPD - as expected, so this was good.
    • Then, with the PRMI locked, I repeated the test. If there is no DC light field (as expected for the PRMI in this configuration), I wouldn't expect this drive signal to show up in the DCPDs. But in fact, I do. Again, this supports the presence of some (for now unquantified) contrast defect.

While it would seem from these graphs that the RIN of the LO beam at these frequencies is rather high, it is because of the ADC noise. More whitening (to be installed in the coming days) will allow us to get a better estimate, should be ~1e-6 I think.

I was just playing today, still need to setup some more screens, DTT templates etc to do more tests in a convenient way.

Now, I can think about how to commission this setup interferometrically.

Attachment 1: PRMI_RFlock.pdf
PRMI_RFlock.pdf
Attachment 2: PRMI_RFlock_fullscale.pdf
PRMI_RFlock_fullscale.pdf
  15513   Mon Aug 10 16:52:04 2020 gautamUpdateBHDWorkable setup prepared

All the details are in E2000436, and documents linked from there, I think an elog would be much too verbose. In summary, a workable setup consisting of

  • 2 DCPDs interfaced with the realtime CDS system. Note that because this circuit is single-ended, while the AA and ADC are differential receiving, there is an overall gain of 0.5. Explicitly, for the 300 ohm DC transimpedance, the conversion is ~350 cts/mW.
  • A local oscillator beam delivered via fiber that is mode-matched (roughly) with the IFO AS beam.
  • A PZT mounted mirror to control the homodyne phase. The PZT (S320) is an obsolete part and it's hard to find a datasheet for it, but if its specs are comparable to the more modern S330, the full stroke is 10 um, for a max applied voltage of 100 V DC, so 100nm/V. c.f. 200V for 3um full stroke of the Noliac.

was prepared.

Last night, I locked the PRMI with the carrier resonant, and convinced myself that the DCPD null stream was sensing the MICH degree of freedom (while it was locked on AS55_Q) with good SNR below ~60 Hz. Above ~60 Hz, in this configuration, the ADC noise was dominating, but by next week, I'll have a whitening board installed that will solve this particular issue. With the optical gain of MICH in this configuration, the ADC noise level was equivalent to ~500 nrad/rtHz of phase noise above ~60 Hz (plots later).

Now, I can think about how to commission this setup interferometrically.

  15512   Mon Aug 10 07:13:00 2020 YehonathanUpdateBHDMonte Carlo Simulations

I fixed some stuff in the MCMC simulation:

1. Results are now plotted as shades from minimum to maximum. I tried making the shade the STD around a mean but it doesn't look good on a log scale when the STD is bigger than the mean.

2. Added comparison with aLigo. The OMCL diff and comm motions in A+ are both compared to the single OMCL DOF of aLigo.

3. I fixed a serious error in the code that produced incorrect results.

4. Imbalances in the IFO such as differential arm loss are generated randomly at the beginning and stay fixed for the rest of the simulation instead of being treated as an offset.

5. The simulation now runs with maxtem=2. That is, TEM modes up to 2nd order are considered.

The results are attached.

 

Attachment 1: MICH_AplusMCMC.pdf
MICH_AplusMCMC.pdf
Attachment 2: PRCL_AplusMCMC.pdf
PRCL_AplusMCMC.pdf
Attachment 3: SRCL_AplusMCMC.pdf
SRCL_AplusMCMC.pdf
Attachment 4: OMC_Comm_AplusMCMC.pdf
OMC_Comm_AplusMCMC.pdf
Attachment 5: OMC_Diff_AplusMCMC.pdf
OMC_Diff_AplusMCMC.pdf
Attachment 6: OMC_Angle_Yaw_AplusMCMC.pdf
OMC_Angle_Yaw_AplusMCMC.pdf
Attachment 7: OMC_Angle_Pitch_AplusMCMC.pdf
OMC_Angle_Pitch_AplusMCMC.pdf
Attachment 8: L0_RIN_AplusMCMC.pdf
L0_RIN_AplusMCMC.pdf
  15510   Sat Aug 8 07:36:52 2020 Sanika KhadkikarConfigurationCalibration-RepairBS Seismometer - Multi-channel calibration

Summary : 

I have been working on analyzing the seismic data obtained from the 3 seismometers present in the lab. I noticed while looking at the combined time series and the gain plots of the 3 seismometers that there is some error in the calibration of the BS seismometer. The EX and the EY seismometers seem to be well-calibrated as opposed to the BS seismometer.

The calibration factors have been determined to be :

BS-X Channel: \small {\color{Blue} 2.030 \pm 0.079 }

BS-Y Channel: \small {\color{Blue} 2.840 \pm 0.177 }

BS-Z Channel: \small {\color{Blue} 1.397 \pm 0.182 }


Details :

The seismometers each have 3 channels i.e X, Y, and Z for measuring the displacements in all the 3 directions. The X channels of the three seismometers should more or less be coherent in the absence of any seismic excitation with the gain amongst all the similar channels being 1. So is the case with the Y and Z channels. After analyzing multiple datasets, it was observed that the values of all the three channels of the BS seismometer differed very significantly from their corresponding channels in the EX and the EY seismometers and they were not calibrated in the region that they were found to be coherent as well. 


Method :

Note: All the frequency domain plots that have been calculated are for a sampling rate of 32 Hz. The plots were found to be extremely coherent in a certain frequency range i.e ~0.1 Hz to 2 Hz so this frequency range is used to understand the relative calibration errors. The spread around the function is because of the error caused by coherence values differing from unity and the averages performed for the Welch function. 9 averages have been performed for the following analysis keeping in mind the needed frequency resolution(~0.01Hz) and the accuracy of the power calculated at every frequency. 

  1. I first analyzed the regions in which the similar channels were found to be coherent to have a proper gain analysis. The EY seismometer was found to be the most stable one so it has been used as a reference. I saw the coherence between similar channels of the 2 seismometers and the bode plots together. A transfer function estimator was used to analyze the relative calibration in between all 3 pairs of seismometers. In the given frequency range EX and EY have a gain of 1 so their relative calibration is proper. The relative calibration in between the BS and the EY seismometers is not proper as the resultant gain is not 1. The attached plots show the discrepancies clearly : 
  • BS-X & EY-X Transfer Function : Attachment #1
  • BS-Y & EY-Y Transfer Function : Attachment #2

          The gain in the given frequency range is ~3. The phase plotting also shows a 180-degree phase as opposed to 0 so a negative sign would also be required in the calibration factor. Thus the calibration factor for the Y channel of the BS seismometer should be around ~3. 

  • BS-Z & EY-Z Transfer Function : Attachment #3

The mean value of the gain in the given frequency range is the desired calibration factor and the error would be the mean of the error for the gain dataset chosen which is caused due to factors mentioned above.

Note: The standard error envelope plotted in the attached graphs is calculated as follows :

         1. Divide the data into n segments according to the resolution wanted for the Welch averaging to be performed later. 

         2. Calculate PSD for every segment (no averaging).

         3. Calculate the standard error for every value in the data segment by looking at distribution formed by the n number values we obtain by taking that respective value from every segment.

Discussions :

The BS seismometer is a different model than the EX and the EY seismometers which might be a major cause as to why we need special calibration for the BS seismometer while EX and EY are fine. The sign flip in the BS-Y seismometer may cause a lot of errors in future data acquisitions. The time series plots in Attachment #4 shows an evident DC offset present in the data. All of the information mentioned above indicates that there is some electrical or mechanical defect present in the seismometer and may require a reset. Kindly let me know if and when the seismometer is reset so that I can calibrate it again. 

Attachment 1: BS_X-EY_X.png
BS_X-EY_X.png
Attachment 2: BS_Y-EY_Y.png
BS_Y-EY_Y.png
Attachment 3: BS_Z-EY_Z.png
BS_Z-EY_Z.png
Attachment 4: timeseries.png
timeseries.png
  15509   Fri Aug 7 11:23:47 2020 ranaUpdateBHDMonte Carlo Simulations

that's great. I think we would like to figure out how to present this so that its clear what the distribution of TFs is. Maybe we can plot the most likely curve as well as a shaded region indicating the 5% and 95% values?

Quote:

I've pushed an MCMC simulation to the A+ BHD repo (filename MCMC_TFs.ipynb). The idea is to show how random offsets around ideal IFO change the noise couplings of different DOFs to readout.

and then we add the loops

  15508   Thu Aug 6 22:57:20 2020 gautamUpdatesafetyNew live HV Supplies

Be aware that there is now a KEPCO HV supply that is energized, sitting on the floor immediately adjacent to the OMC rack, east of the AP table. It is currently set to 100 V DC, and a PI PZT installed on the AP table has its 3 PZTs energized by said supply (via an OMC piezo driver). I will post pictures etc of the work from the last 10 days over the weekend.

  15507   Thu Aug 6 00:34:38 2020 YehonathanUpdateBHDMonte Carlo Simulations

I've pushed an MCMC simulation to the A+ BHD repo (filename MCMC_TFs.ipynb). The idea is to show how random offsets around ideal IFO change the noise couplings of different DOFs to readout.

At each step of the simulation:

1. Random offsets for the different DOFs are generated from a normal distribution. The RMSs are taken from experimental data and some guesses and can be changed later. The laser frequency is tuned to match the CARM offset.

These are the current RMS detunings I use:

DOF RMS Taken from
DARM 10fm PHYSICAL REVIEW D 93, 112004 (2016), Table 2
CARM 1fm PHYSICAL REVIEW D 93, 112004 (2016), Table 2
MICH 3pm PHYSICAL REVIEW D 93, 112004 (2016), Table 2
PRCL 1pm PHYSICAL REVIEW D 93, 112004 (2016), Table 2
SRCL 10pm PHYSICAL REVIEW D 93, 112004 (2016), Table 2
OMCL 0.1pm Guess
OMC Breadboard angle 1\mu rad Guess
Differential arm loss 15ppm Guess
BHD BS imbalance 10% Guess
OMC finesse imbalance 5ppm Guess

2. A transfer function is computed for the noisy DOFs.

3. Projected noise is calculated.

These are the noise level for the DOFs:

DOF Noise Taken from
MICH 2e-16 m PHYSICAL REVIEW D 93, 112004 (2016), Fig 9
PRCL 0.5e-17 m PHYSICAL REVIEW D 93, 112004 (2016), Fig 9
SRCL 5e-16 PHYSICAL REVIEW D 93, 112004 (2016), Fig 9
OMCL 2.5e-17*(100/f)^(1/2) LIGO-G1800149
OMC Breadboard angle 1nrad Guess
RIN 2e-9 Optics Letters Vol. 34, Issue 19, pp. 2912-2914 (2009)

 

The attachments show the projected noise levels for the noisy DOFs. Each curve is a different instance of random offsets. The ideal case - "zero offsets" is also shown.

OMC Comm and OMC diff refer to the common and differential length change of the OMCs.

Attachment 1: MICH_Aplus_MCMC.pdf
MICH_Aplus_MCMC.pdf
Attachment 2: PRCL_Aplus_MCMC.pdf
PRCL_Aplus_MCMC.pdf
Attachment 3: SRCL_Aplus_MCMC.pdf
SRCL_Aplus_MCMC.pdf
Attachment 4: OMC_Comm_Aplus_MCMC.pdf
OMC_Comm_Aplus_MCMC.pdf
Attachment 5: OMC_Diff_Aplus_MCMC.pdf
OMC_Diff_Aplus_MCMC.pdf
Attachment 6: OMC_Angle_Yaw_Aplus_MCMC.pdf
OMC_Angle_Yaw_Aplus_MCMC.pdf
Attachment 7: OMC_Angle_Pitch_Aplus_MCMC.pdf
OMC_Angle_Pitch_Aplus_MCMC.pdf
Attachment 8: L0_RIN_Aplus_MCMC.pdf
L0_RIN_Aplus_MCMC.pdf
  15506   Thu Jul 30 16:16:43 2020 gautamUpdateSUSSuspension recovery

This earthquake and friends had tripped all watchdogs. I used the scripted watchdog re-enabler, and released the stuck ITMX (this operation is still requires a human and hasn't been scripted yet). IMC is locked again and all Oplevs report healthy optic alignment.

  15505   Wed Jul 29 11:57:59 2020 ranaUpdateBHDIn-air BHD - CDS and wiring summary

3. I agree - this whitening will be handy to have for diagnostics.

4. I think in principle, we can ask a company to make the custom cables for us to save us some hand labor. Rich/Chub probably know the right companies to do small numbers of dirty cables.

5. Can't we just a single Noliac PZT in the same way that the OMC does? Or is the lead time too long?

6. Do we need active steering for this in-air test? I'm not even sure how we would get the alignment signal, so maybe that's a good reason to figure this out.

  15504   Tue Jul 28 14:11:14 2020 JonUpdateVACVac interlock test today 1:30 pm

This test has been completed. The IFO configuration has been reverted to nominal.

For future reference: yes, both the V4 and V5 hardware interlocks were found to still be connected and work. A TTL signal from the analog output port of each pump controller (TP2 and TP3) is connected to an auxiliary relay inside the main valve relay box. These serve the purpose of interupting the (Acromag) control signal to the primary V4/5 relay. This interrupt is triggered by each pump's R1 setpoint signal, which is programmed to go low when the rotation speed falls below 80% of the low-speed setting.

Quote:

This afternoon Jordan is going to carry out a test of the V4 and V5 hardware interlocks. To inform the interlock improvement plan [15499], we need to characterize exactly how these work (they pre-date the 2018 upgrade). I have provided him a sequence of steps for each test and will also be backing him up on Zoom.

We will close V1 as a precaution but there should be no other impact to the IFO. The tests are expected to take <1 hour. We will advise when they are completed.

  15503   Tue Jul 28 13:55:11 2020 HangUpdateBHDExploring bilinear SRCL->DARM coupling

We explore bilinear SRCL to DARM noise coupling mechanisms, and show two cases that by doing BHD readout the noise performance can be improved. In the first case, the bilinear piece is due to residual DHARD motion (see also LHO:45823), and it matters mostly for the low-frequency (<100 Hz) part, and in the second piece the bilinear piece is due to residual SRCL fluctuation and it matters mostly for the a few x 100 Hz part. Details are below:

=================================================

General Model:

We can write the SRCL to DARM transfer function as (Evan Hall's thesis, eq. 2.29)

Z_s2d(f) = C_lf(f) * F^2 * x_D + C_hf(f) * F * dphi_S * x_D    ---- (1)

where

C_lf ~ 1/f^2 and C_hf ~ f are constants at each frequency unless there are major upgrades to the IFO,

F is the finesse of the arm cavity which depends on the alignment, spot position on the TMs, etc., 

dphi_S is the SRCL detuning (wrt the nominal 90 deg value), 

x_D is the DC DARM offset. 

The linear part of this can be removed with feedforward subtractions and it is the bilinear piece that matters, which reads

dZ_s2d = C_lf * <F>^2 * dx_D + C_hf * <F> * <dphi_S> * dx_D

             + 2C_lf * <F> * <x_D>  * dF + C_hf * <dphi_S> * <x_D> * dF

             + C_hf  * <F> * <x_D> * d(dphi_S).     ---- (2)

The first term in (2) is due to residual DARM motion dx_D. This term does not depends on the DC value of DARM offset <x_D> and thus does not depend on doing BHD or DC readout. On the other hand, the typical residual DARM motion is 1 fm << 1 pm of DARM offset. Since the current feedforward reduction factor is about 10 (see both Den Martynov's thesis and Evan Hall's thesis), clearly we are not limited by the residual DARM motion. 

The second term is due to the change in the arm finesse, which can be affected by, e.g., the alignment fluctuation (both increasing the loss due to scattering into 01/10 modes and affecting the spot positon and hence changing the losses), and is likely to be the reason why we see the effect being modulated by DHARD. 

The last term in (2) is due to the residual SRCL fluctuation and is important for the ~ a few x 100 Hz band.

=================================================

DHARD effects. 

As argued above, the DHARD affects the SRCL -> DARM coupling as it changes the finesse in the arm cavity (through scattering into 01/10 modes; in finesse we cannot directly simulate the effects due to spot hitting a rougher location). 

Since in the second term of eq. (2) the LF part depends on the DARM DC offset <x_D>, this effect can be improved by going from DC readout to BHD. 

To simulate it in finesse, at a fixed DARM DC offset, we compute the SRCL->DARM transfer functions at different DHARD offsets, and then numerically compute the derivative \partial Z_s2d / \partial \theta_{DH}. Then multiplying this derivative with the rms value of DHARD fluctuation \theta_{DH} we then know the expected bilinear coupling piece. 

The result is shown in the first attached plot. Here we have assumed a flat SRCL noise of 5e-16 m/rtHz for simplicity (see PRD 93, 112004, 2016). We do not account for the loop effects which further reduces the high frequency components for now. The residual DHARD RMS is assumed to be 1 nrad. 

In the first plot, from top to bottom we show the SRCL noise projection at different DARM DC offsets of (0.1, 1, 10) pm. Since the DHARD alignment only affects the arm finesse starting at quadratic order, it thus matters what DC offset in DHARD we assume. In each pannel, the blue trace is for no DC offset in DHARD and the orange one for a 5 nrad DC offset. As a reference, the A+ sensitivity is shown in grey trace in each plot as a reference. 

We can see if there is a large DC offset in DHARD (a few nrad) and we still do DC readout with a few pm of DARM offset, then the bilinear piece of SRCL can still contaminate the sensitivity in the 10-100 Hz band (bottom panel; orange trace). On the other hand, if we do BHD, then the SRCL noise should be down by ~ x100  even compared to with the top panel. 

(A 5 nrad of DC offset in DHARD coupled with 1 nrad RMS would cause about 0.5% RIN in the arms. This is somewhat greater than the typically measured RIN which is more like <~ 0.2%. See the second plot). 

=================================================

SRCL effect. 

Similarly we can consider the SRCL->DARM coupling due to residual SRCL rms. The approach is very similar to what we did above for DHARD. I.e., we compute Z_s2d at fixed DARM offset and for different SRCL offsets, then we numerically evaluate \partial Z_s2d / \partial dphi_S. A residual SRCL rms of 0.1 nm is then used to generate the projection shown in the third figure. 

Unlike the DHARD effect, the bilinear SRCL piece does not depend on the DC SRCL detuning (for the 50-500 Hz part). It does still depends on the DARM DC offset and therefore could be improved by BHD.

Since we do not include the LP of the SRCL loop in this plot, the HF noise at 1 kHz is artifical as it can be easily filtered out. However, the LP will not be very strong around 100-300 Hz for a SRCL UGF ~ 30 Hz, and thus doing BHD could still have some small improvements for this effect. 

Attachment 1: SRCL_bilin_DHARD.pdf
SRCL_bilin_DHARD.pdf
Attachment 2: ARM_RIN.pdf
ARM_RIN.pdf
Attachment 3: SRCL_bilin_SRCL.pdf
SRCL_bilin_SRCL.pdf
  15502   Tue Jul 28 12:22:40 2020 JonUpdateVACVac interlock test today 1:30 pm

This afternoon Jordan is going to carry out a test of the V4 and V5 hardware interlocks. To inform the interlock improvement plan [15499], we need to characterize exactly how these work (they pre-date the 2018 upgrade). I have provided him a sequence of steps for each test and will also be backing him up on Zoom.

We will close V1 as a precaution but there should be no other impact to the IFO. The tests are expected to take <1 hour. We will advise when they are completed.

  15501   Mon Jul 27 15:48:36 2020 JonSummaryVACVacuum parts ordered

To carry out the next steps of the vac refurbishment plan [ELOG 15499], I've ordered parts necessary for interfacing the UPS units and the analog TP2/3 controller outputs with c1vac. The purchase list is appended to the main BHD list and is located here. Some parts we already had in the boxes of Acromag materials. Jordan is gathering what we do already have and staging it on the vacuum controls console table - please don't move them or put them away.

Quote:

Replace failing UPS.

Remove interlock dependencies on TP2/TP3 serial readbacks. Due to persistent glitching [ELOG 15140, ELOG 15392].

  15500   Fri Jul 24 15:40:59 2020 JordanUpdateVACInstallation of two new UPS units

I installed the Tripp Lite SMX1000RT2U and Tripp Lite Smart1000LCD at the bottom of the 1x8 electronics rack. These are plugged in to power, and are ready for testing. All other cables (serial, usb, etc.) have been left on the table next to the 1x8 rack.

Attachment 1: UPS.jpg
UPS.jpg
  15499   Thu Jul 23 15:58:24 2020 JonSummaryVACVacuum controls refurbishment plan

This year we've struggled with vacuum controls unreliability (e.g., spurious interlock triggers) caused by decaying hardware. Here are details of the vacuum refurbishment plan I described on the 40m call this week.

 Refurbish TP2 and TP3 dry pumps. Completed [ELOG 15417].

 Automated notifications of interlock-trigger events. Email to 40m list and a new interlock flag channel. Completed [ELOG 15424].

Replace failing UPS.

  • Two new Tripp Lite units on order, 110V and 230V [ELOG 15465].
  • Jordan will install them in the vacuum rack once received.
  • Once installed, Jon will come test the new units, set up communications, and integrate them into the interlock system following this plan [ELOG 15446].
  • Jon will move the pumps and other equipment to the new UPS units only after completing the above step.

Remove interlock dependencies on TP2/TP3 serial readbacks. Due to persistent glitching [ELOG 15140, ELOG 15392].

Unlike TP2 and TP3, the TP1 readbacks are real analog signals routed to Acromags. As these have caused us no issues at all, the plan is to eliminate dependence on the TP2/3 digital readbacks in favor of the analog controller outputs. All the digital readback channels will continue to exist, but the interlock system will no longer depend on them. This will require adding 2 new sinking BI channels each for TP2 and TP3 (for a total of 4 new channels). We have 8 open Acromag XT1111 channels in the c1vac system [ELOG 14493], so the new channels can be accommodated. The below table summarizes the proposed changes.

Channel Type Status Description Interlock
C1:Vac-TP1_current AI exists Current draw (A) keep
C1:Vac-TP1_fail BI exists Critical fault has occurred keep
C1:Vac-TP1_norm BI exists Rotation speed is within +/-10% of set point new
C1:Vac-TP2_rot soft exists Rotation speed (krpm) remove
C1:Vac-TP2_temp soft exists Temperature (C) remove
C1:Vac-TP2_current soft exists Current draw (A) remove
C1:Vac-TP2_fail BI new Critical fault has occurred new
C1:Vac-TP2_norm BI new Rotation speed is >80% of set point new
C1:Vac-TP3_rot soft exists Rotation speed (krpm) remove
C1:Vac-TP3_temp soft exists Temperature (C) remove
C1:Vac-TP3_current soft exists Current draw (A) remove
C1:Vac-TP3_fail BI new Critical fault has occurred new
C1:Vac-TP3_norm BI new Rotation speed is >80% of set point new
  15498   Tue Jul 21 16:41:46 2020 gautamUpdateBHDPMC assembly space

I decided to use the old EY auxiliary optics table, which is now stored along the east arm about 10 m from the end, as a workspace for assembling the little PMCs. I wiped everything down with isopropanol for general cleanliness, removed the metal plate on the south edge of the table enclosure to allow access, covered the table with some clean Aluminium foil, and then moved the plastic box with PMC parts to the table - see Attachment #1. I haven't actually done any assembly just yet, waiting for more info (if available) on the procedure and implements available...

Attachment 1: IMG_8635.JPG
IMG_8635.JPG
  15497   Tue Jul 21 00:30:24 2020 gautamUpdateBHDIn-air BHD - LO RIN

Attachment #1 shows the RIN of the local oscillator beam delivered to the AP table via fiber. I used a PDA520 to make this measurement, while the electronics for the DCPDs are pending. I don't really have an explanation for the difference between the locked IFO trace vs the not locked trace - we don't have an ISS running (but this first test suggests we should) and the beam is picked off before any cavities etc, so this is a reflection of the state of the FSS servo at the times of measurement?


Tried locking CARM using the hybrid REFL (for AO path) and POX 11 (for MCL path) scheme a bunch of times today, but I had no luck. When the CARM offset is zeroed, the PRMI lock is lost almost immediately. Maybe this is indicative of some excess noise in the POX data stream relative to the REFL signal? The one thing I haven't tried is to take the IFO all the way to the locked state, and then transition the MCL actuation from CM_SLOW to POX11_I.


An SR785 is sitting on the North side of the AP table in the walkway - I will clear it tomorrow.

Attachment 1: LO_RIN.pdf
LO_RIN.pdf
ELOG V3.1.3-